# Spin waves and three-dimensionality in the high-pressure antiferromagnetic phase of SrCu2(BO3)2

### v1

@@ Abstract
SrCu2(BO3)2 is an ideally frustrated system composed of Cu2+ dimers ($S=1/2$ spin pairs) arranged orthogonally both in and between quasi-two-dimensional (2D) layers, as represented in Figs. 1(a)-(c) Kageyama et al. (1999a). The in-plane network realizes the Shastry-Sutherland model (SSM) [inset Fig. 1(d)], proposed Shastry and Sutherland (1981) for its exact dimer-singlet ground state at small $J^{\prime}/J$ and governed by this ratio of competing interactions Koga and Kawakami (2000); Chung et al. (2001); Läuchli et al. (2002); Corboz and Mila (2013); Nakano and Sakai (2018); Xi et al. (2023). SrCu2(BO3)2 exhibits the dimer-singlet ground state at ambient pressure Miyahara and Ueda (1999); Kageyama et al. (1999a, 2000a, 2000b); Gaulin et al. (2004); Kakurai et al. (2005) and is well described by a SSM with $J^{\prime}/J=0.63$ and additional weak (3%) Dzyaloshinskii-Moriya (DM) interactions Cépas et al. (2001); Miyahara and Ueda (2003); Kodama et al. (2005); Miyahara and Mila (2005); Shi et al. (2022). Remarkably, the phase diagram of SrCu2(BO3)2 under an applied hydrostatic pressure Waki et al. (2007); Zayed et al. (2017); Sakurai et al. (2018), shown in Fig. 1(d), mirrors that of the SSM with increasing $J^{\prime}/J$ [Fig. 1(e)], and is studied extensively for its exotic QPTs Guo et al. (2020); Larrea Jiménez et al. (2021); Guo et al. (2023). Above 4 GPa, where the tetragonal symmetry changes to monoclinic Loa et al. (2005); Haravifard et al. (2014); Zayed et al. (2014), antiferromagnetic (AFM) order is found by neutron and X-ray diffraction Haravifard et al. (2014); Zayed et al. (2017); Zayed (2010) and phase-transition features are reported around 8 K Guo et al. (2020) and above 120 K Haravifard et al. (2014); Zayed et al. (2014); Guo et al. (2020). However, the high pressures pose a severe challenge to experiment and rather little is known about either the precise pressure regimes for the SSM-AFM and monoclinic-AFM phases or the possible magnetic interactions in either phase.
@@
Our INS experiments used a 58 mg disc-shaped single-crystalline sample of SrCu2(BO3)2, which was grown by the floating-zone method Kageyama et al. (1999b); Jorge et al. (2004) and oriented with ${\bf Q}=(q_{h},q_{k},0)$ in the horizontal scattering plane. The small sample size is a further challenge for INS, but recent developments in high-pressure techniques at the Institute Laue-Langevin (ILL) Komatsu et al. (2014); Klotz et al. (2019); Funnell et al. (2021) made this experiment possible. Using a Paris-Edinburgh press, the pressure was increased gradually to 5.5 GPa while the lattice parameters of SrCu2(BO3)2 Loa et al. (2005) were measured for pressure determination. The sample chamber was pre-cooled with liquid N2 and cooling to 4.5 K was provided by a closed-cycle refrigerator. We performed two consecutive experiments at the IN8 and ThALES triple-axis spectrometers with the same loading. We used fixed momentum transfers $k_{f}=2.66$-1 at IN8 and $k_{f}=1.50$-1 at ThALES, providing respective instrumental resolutions (FWHM) of $1.05(1)$ and $0.167(4)$ meV. Scans with constant energy transfer, $E$, or constant momentum transfer, $\bf Q$, were performed with typical counting times of 5 min per point. In the Supplemental Material (SM) sup we provide detailed descriptions of the pressure loading procedure (Sec. S1A), instrument configurations (Sec. S1B), pressure determination (Sec. S1C), and data treatment (Sec. S1D). Supporting diffraction data were collected at IN3, also at the ILL, and this experiment is described in Sec. S1E.
@@
A selection of constant-$E$ and constant-$\bf Q$ scans is shown in Fig. 2, with data from both IN8 and ThALES. The width of the elastic line on IN8 allowed for access to energy transfers above approximately 2 meV, whereas at ThALES 0.5 meV was accessible. Because the atomic positions in the monoclinic high-pressure phase remain close to those of the tetragonal phase, with in-plane lattice parameters $a/b\simeq 0.999$ (and monoclinic angle $\beta\simeq 94^{\circ}$) Haravifard et al. (2014), we retain the tetragonal approximation with $a^{*}$ and $b^{*}=2\pi/a$. Figures 3(a)-(b) collect all neutron intensities measured respectively along $(0,q_{k},0)$ and $(q_{h},1,0)$ at constant $E$, and make clear that dispersive modes are present with origins at $(0,1,0)$ and $(1,1,0)$. Above 5 meV it becomes difficult to track the mode positions due to vanishing neutron intensity [Figs. 2(a) and 2(d)]. At 1.5 meV and below, we no longer observe neutron intensity at $(0,1,0)$ [Figs. 2(b) and 2(c)]. In Figs. 2(a)-(b) we fitted the data collected along $(q_{h},1,0)$ with Gaussian line shapes centered respectively at ±$\pm\delta_{h}$ and ±$\pm\delta_{h}^{\prime}$ from $(0,1,0)$ and $(1,1,0)$, allowing the peak widths and intensities to vary individually (Sec. 1D of the SM sup ). Similar fits to the data collected along $(0,q_{k},0)$ yielded the mode positions along both directions in $\bf Q$. The dispersion minimum at $(0,1,0)$ is resolved in a constant-$\bf Q$ scan from ThALES [Fig. 2(c)]. A Gaussian fit yields an energy of 1.90(8) meV and the magnetic origin of the excitation is demonstrated by its absence at 140 K. In Fig. 2(d), we show constant-$\bf Q$ scans from IN8 where subtraction of the incoherent contribution allowed the fitting of one or two Gaussian profiles at different $\bf Q$. The fitted mode positions are also marked in Figs. 3(a)-(b).
@@
Figure 2: Selection of INS data collected at IN8 and ThALES at 5.5 GPa. (a)-(b) Constant-$E$ scans along ${\bf Q}=(q_{h},1,0)$. (c)-(d) Constant-$\bf Q$ scans. Fits to determine the mode positions in $\bf Q$ and $E$ are illustrated with solid lines and the shaded areas represent the individual Gaussians. Dashed lines in panel (d) show the incoherent contribution to the scattering signal measured across the elastic line at $(0.7,1.3,0)$.
@@
Figure 3: Measured and modelled excitation spectrum of SrCu2(BO3)2 in the monoclinic-AFM phase. (a)-(b) INS intensity collected on IN8 at 5.5 GPa and 4.5 K, shown as a function of $E$ and $\bf Q$ along the $(0,q_{k},0)$ and $(q_{h},1,0)$ directions. The inset in panel (b) shows lower-energy data (dashed box) collected at ThALES around $(0,1,0)$. (c)-(d) Comparison with the calculated spectrum using an instrumental energy resolution of 1 meV. The splittings, $\Delta$ and $\Delta^{\prime}$, are indicated in panel (c). The mode positions determined from the INS data are shown with open pink circles (IN8) and open pink triangles (ThALES). The orientation of the error bars indicates whether a given data point originates from a constant-$E$ or constant-$\bf Q$ scan. The grey curves show the dispersion calculated from linear spin-wave theory.
@@
From the near-tetragonal nature of the high-pressure phase, we assume that the SSM continues to provide a minimal model for the observed spin-wave dispersion. The physics of the AFM phase is that the stronger $J^{\prime}$ forces the spins to order ferromagnetically (FM) on the dimers and AFM between dimers, both in and out of plane [Fig. 1(c)]. The 2D AFM phase has two doubly degenerate spin-wave branches, a Goldstone mode and an optical mode. Although we have no dispersion measurements in the $c$ direction, the observed in-plane scattering intensity [Figs. 3(a)-(b)] motivates the introduction of an inter-layer interaction, $J_{c}$, between orthogonal dimers along the $(0,0,1)$ direction, as shown in Fig. 1(c). This model has the following properties: (i) the degeneracy of the low-energy mode is lifted, with a splitting $\Delta=8S\sqrt{J^{\prime}J_{c}}$ above the massless point; (ii) due to the AFM inter-layer stacking, the intensity of the Goldstone mode vanishes at ${\bf Q}=(0,1,0)$ and $(0,3,0)$, making the spectrum look gapped at these positions; (iii) there is an intensity asymmetry along $(0,q_{k},0)$ around $q_{k}=1$ and $3$; (iv) the splitting between the low-energy and optical branches at the top of the former [e.g. ${\bf Q}=(0,1.5,0)$ and $(0.5,1,0)$] is $\Delta^{\prime}\approx 2SJ$.
@@
We refine the model parameters based on the experimentally determined mode positions and intensities with assistance from spinW Toth and Lake (2015), a software implementation of linear spin-wave theory, to obtain $J^{\prime}=4.28(14)$ meV, $J=2.36(26)$ meV (i.e. $J^{\prime}/J=1.8(2)$), and $J_{c}=0.053(3)$ meV. The spectra for an assumed instrumental energy resolution of 1 meV are shown in Figs. 3(c)-(d), and provide good agreement with the measurements in Figs. 3(a)-(b). In particular, the fitted spectrum confirms an AFM layer stacking, because FM stacking would result in Goldstone-mode intensity at $(0,1,0)$ and $(0,3,0)$, but not at $(1,1,0)$. We note that our deduced parameter values are not corrected for the quantum renormalization factor, $Z_{c}$, which to our knowledge has not been evaluated for the AFM phase of the SSM.
@@
Currently the most refined estimates for the interaction parameters of SrCu2(BO3)2 at ambient pressure are $J=81.5$ K (7.0 meV) and $J^{\prime}=51.3$ K (4.4 meV, $J^{\prime}/J=0.63$) Shi et al. (2022). While there is no exact model for the increase of $J^{\prime}/J$ with pressure, a linear decrease of both $J$ and $J^{\prime}$, with $\Delta J/P$ falling 3-4 times faster than $\Delta J^{\prime}/P$ Zayed et al. (2017), reproduces the $(H,P)$ phase diagram rather well in the dimer and plaquette phases Shi et al. (2022). Ab initio calculations predict a discontinuity in the parameters on entering the SSM-AFM phase Badrtdinov et al. (2020), and at higher pressures no information for $J$ or $J^{\prime}$ was available until now. Although it is unexpected that $J$ is suppressed so strongly, it is not implausible. Its size and sign are highly sensitive to the Cu-O-Cu bond angle following the Goodenough-Kanamori rules Goodenough (1955); Kanamori (1959), especially when this angle is close to 95 Mizuno et al. (1998). Measurements in SrCu2(BO3)2 [Figs. 1(a)-(b)] indicate a decrease from $98^{\circ}$ at ambient pressure to $94^{\circ}$ at 3.7 GPa Zayed (2010). The $J^{\prime}$ superexchange path is through a BO3 unit, which remains robust with pressure and locked in position by a mirror plane, becoming free to rotate out of the plane only beyond the monoclinic transition. Our measurements therefore quantify all of these considerations.
@@
Turning to $J_{c}$, we note first that the system is much softer along the ${\hat{c}}$ direction than in the plane, with the $c$ lattice parameter reduced by around 6% at 5.5 GPa, whereas $a$ and $b$ are reduced by less than 1% Loa et al. (2005); Haravifard et al. (2014). The buckling of the Cu-BO3 planes shown in Fig. 1(b) leads to two independent $J_{c}$ parameters Miyahara and Ueda (2003). However, the in-plane INS spectrum is sensitive only to their combined effect, and thus we fit a single effective $J_{c}$; further details are presented in Sec. S2A of the SM sup . In this quasi-2D system, the splitting $\Delta$ is a steep function of $J_{c}$ and in our INS experiment $\Delta=1.90(8)$ meV is measured to high accuracy, making our estimate of $J_{c}=0.053(3)$ meV very precise. Its small value reflects the absence of significant inter-layer superexchange paths, and is likely to be smaller still at ambient pressure. An indirect estimate of the ambient-pressure $J_{c}$ as a 10% correction to the high-temperature magnetic susceptibility Miyahara and Ueda (2000, 2003) returned a value of 0.7 meV, an order-of-magnitude difference from our result that is likely a consequence of strong finite-size effects inherent to early numerical studies (since remedied by more modern technology Wietek et al. (2019)).
@@
Although a Heisenberg SSM with an inter-layer coupling captures all the features of the spin-wave dispersion in the high-pressure phase, SrCu2(BO3)2 is known to have weak DM interactions at ambient pressure Miyahara and Ueda (2003); Kodama et al. (2005); Shi et al. (2022). These terms also act to gap the Goldstone mode, but the DM interaction required to create a gap of 1.9 meV in a 2D SSM is comparable to $J$, rendering this an unlikely mechanism to explain the INS spectra. Although our minimal model has no DM interactions, we absolutely do not claim that they are no longer present, only that we cannot resolve them. In Sec. 2B of the SM sup we explore models including DM interactions, and show that even 10% DM terms in the lower-symmetry monoclinic phase of SrCu2(BO3)2 provide only a minor correction to the Heisenberg interaction parameters.
@@
The confirmation of a small $J_{c}$ is essential information for assessing the suitability of a 2D SSM to describe SrCu2(BO3)2. The fact that the dimer ($J$) bonds are forced into a FM configuration in the AFM phase makes the four $J_{c}$ bonds in an inter-layer tetrahedron fully unfrustrated [Fig. 1(c)], and hence they contribute an energy of order 1 K per site to the stability of a 3D AFM phase. By contrast, in the dimer-singlet phase these bonds are simply ineffective. Concerning the plaquette phase, it was shown rather early that its very existence depends on having only a weak $J_{c}$ Koga (2000), and the propensity of $J_{c}$ to destabilize the plaquette phase has been used recently to extract a value $J_{c}\approx 0.2$ meV Vlaar and Corboz (2023). Our definitive $J_{c}$ can be used to refine the estimates of intra-layer energetics and their pressure evolution on which that estimate was based. Another delicate issue is whether the plaquette singlets form around the squares that contain the dimer bond (“full plaquette”) Zayed et al. (2017) or the squares that do not (“empty plaquette”) Corboz and Mila (2013). The minimal $J_{c}$ we consider is thought not to affect plaquette formation Vlaar and Corboz (2023), and rather too little is known about the 3D plaquette order to make a statement on whether more complex inter-layer coupling could do so. Finally, because the plaquette-AFM transition of the 2D SSM has been proposed as a candidate deconfined quantum critical point (DQCP), and SrCu2(BO3)2 as a candidate material for its observation Lee et al. (2019); Yang et al. (2022); Cui et al. (2023), the effects of inter-layer coupling should definitely be included as part of evaluating such a scenario.
@@
Finally, we remark that our results do not provide an immediate understanding of the reported phase transitions at 8 K Guo et al. (2020) and 120 K Haravifard et al. (2014); Zayed et al. (2014); Guo et al. (2020), the parameters of our quasi-2D model being more likely to favor a single ordering transition at an intermediate temperature (of order $J^{\prime}/2$ Yasuda et al. (2005)). We performed detailed diffraction studies of the $(0,1,0)$ magnetic Bragg peak, as described in Sec. S1E of the SM sup , finding an unconventional thermal evolution with a characteristic temperature of 88 K. While one may speculate about the onset of 2D and of true 3D order, these contradictions underline the need for more systematic combined studies of the structure and magnetism at and above 3 GPa.

### v3

@@ Abstract
SrCu2(BO3)2 is an ideally frustrated system composed of Cu2+ dimers ($S=1/2$ spin pairs) arranged orthogonally both in and between quasi-two-dimensional (2D) layers, as represented in Figs. 1(a)-(c) Kageyama et al. (1999a). The in-plane network realizes the Shastry-Sutherland model (SSM) [inset Fig. 1(d)], proposed Shastry and Sutherland (1981) for its exact dimer-singlet ground state at small $J^{\prime}/J$ and governed by this ratio of competing interactions Koga and Kawakami (2000); Chung et al. (2001); Läuchli et al. (2002); Corboz and Mila (2013); Nakano and Sakai (2018); Xi et al. (2023). SrCu2(BO3)2 exhibits the dimer-singlet ground state at ambient pressure Miyahara and Ueda (1999); Kageyama et al. (1999a, 2000a, 2000b); Gaulin et al. (2004); Kakurai et al. (2005) and is well described by a SSM with $J^{\prime}/J=0.63$ and additional weak (3%) Dzyaloshinskii-Moriya (DM) interactions Cépas et al. (2001); Miyahara and Ueda (2003); Kodama et al. (2005); Miyahara and Mila (2005); Shi et al. (2022). Remarkably, the phase diagram of SrCu2(BO3)2 under an applied hydrostatic pressure Waki et al. (2007); Zayed et al. (2017); Sakurai et al. (2018), shown in Fig. 1(d), mirrors that of the SSM with increasing $J^{\prime}/J$ [Fig. 1(e)], and is studied extensively for its exotic QPTs Guo et al. (2020); Larrea Jiménez et al. (2021); J. Guo et al. (2023). Above 4 GPa, where the tetragonal symmetry changes to monoclinic Loa et al. (2005); Haravifard et al. (2014); Zayed et al. (2014), antiferromagnetic (AFM) order is found by neutron and X-ray diffraction Haravifard et al. (2014); Zayed et al. (2017); Zayed (2010) and phase-transition features are reported around 8 K Guo et al. (2020) and above 120 K Haravifard et al. (2014); Zayed et al. (2014); Guo et al. (2020). However, these high pressures pose a severe challenge to experiment and rather little is known about either the precise pressure regimes for the SSM-AFM and monoclinic-AFM phases or the possible magnetic interactions in either phase.
@@
Our INS experiments used a 58 mg disc-shaped single-crystalline sample of SrCu2(BO3)2, whose growth and characterization we summarize in Sec. S1A of the Supplemental Material (SM) sup . The sample was oriented with ${\bf Q}=(q_{h},q_{k},0)$ in the horizontal scattering plane. Its small size is a challenge for INS, but recent developments in high-pressure techniques at the Institute Laue-Langevin (ILL) Komatsu et al. (2014); Klotz et al. (2019); Funnell et al. (2021) made this experiment possible. Using a Paris-Edinburgh press, the pressure was increased gradually to 5.5 GPa while the lattice parameters of SrCu2(BO3)2 Loa et al. (2005) were measured for pressure determination. The sample chamber was pre-cooled with liquid N2 and cooling to 4.5 K was provided by a closed-cycle refrigerator. We performed two consecutive experiments at the IN8 and ThALES triple-axis spectrometers with the same loading. We used fixed momentum transfers $k_{f}=2.66$-1 at IN8 and $k_{f}=1.50$-1 at ThALES, providing respective instrumental resolutions (FWHM) of $1.05(1)$ and $0.167(4)$ meV. Scans with constant energy transfer, $E$, or constant momentum transfer, $\bf Q$, were performed with typical counting times of 5 min per point. In the SM sup we provide detailed descriptions of the pressure loading procedure (Sec. S1B), instrument configurations (Sec. S1C), pressure determination (Sec. S1D), and data treatment (Sec. S1E). Supporting diffraction data were collected at IN3, also at the ILL, and this experiment is described in Sec. S1F.
@@
A selection of constant-$E$ and constant-$\bf Q$ scans is shown in Fig. 2, with data from both IN8 and ThALES. The width of the elastic line on IN8 allowed for access to energy transfers above approximately 2 meV, whereas at ThALES 0.5 meV was accessible. Because the in-plane lattice parameters of the monoclinic phase satisfy $a/b\simeq 0.999$ (with monoclinic angle $\beta\simeq 94^{\circ}$) Haravifard et al. (2014), we retain the tetragonal approximation and use $a^{*}=b^{*}=2\pi/a$. The color contours in Figs. 3(a)-(b) collect all neutron intensities measured respectively along $(0,q_{k},0)$ and $(q_{h},1,0)$ at constant $E$, and make clear that dispersive modes are present with minima at $(0,1,0)$ and $(1,1,0)$. At 1.5 meV and below, we no longer observe neutron intensity at $(0,1,0)$ [Figs. 2(b) and 2(c)]. Above 5 meV it becomes difficult to discern sharp modes due to vanishing neutron intensity [Figs. 2(a) and 2(d)]. In Figs. 2(a)-(b) we fitted the data collected along $(q_{h},1,0)$ with Gaussian line shapes centered respectively at ±$\pm\delta_{h}$ and ±$\pm\delta_{h}^{\prime}$ from $(0,1,0)$ and $(1,1,0)$, allowing the peak widths and intensities to vary individually (Sec. 1E of the SM sup ). Similar fits to the data collected along $(0,q_{k},0)$ yielded mode positions along both directions in $\bf Q$. The dispersion minimum at $(0,1,0)$ is resolved in a constant-$\bf Q$ scan from ThALES [Fig. 2(c)]. A Gaussian fit yields an energy of 1.90(8) meV and the magnetic origin of the excitation is demonstrated by its absence at 140 K. In Fig. 2(d), we show constant-$\bf Q$ scans from IN8 where subtraction of the incoherent contribution allowed the fitting of one or two Gaussian profiles at different $\bf Q$. The fitted mode positions are marked in Figs. 3(a)-(b), where the relatively large error bars in the high-energy range reflect the difficulty in determining the excitation structure here.
@@
Figure 2: Selection of INS data collected at IN8 and ThALES at 5.5 GPa. (a)-(b) Constant-$E$ scans along ${\bf Q}=(q_{h},1,0)$. (c)-(d) Constant-$\bf Q$ scans. Fits to determine the mode positions in $\bf Q$ and $E$ are illustrated with solid lines and the shaded areas represent the individual Gaussians. Dashed lines in panel (d) show the incoherent scattering contribution measured across the elastic line at $(0.7,1.3,0)$. Orange symbols mark the predicted mode centers obtained in a global parameter fit.
@@
Figure 3: Measured and modeled excitation spectra of SrCu2(BO3)2 in the monoclinic-AFM phase. (a)-(b) INS intensity collected on IN8 at 5.5 GPa and 4.5 K, shown as a function of $E$ and $\bf Q$ along $(0,q_{k},0)$ and $(q_{h},1,0)$. The inset in panel (b) shows lower-energy data (dashed box) collected at ThALES around $(0,1,0)$. (c)-(d) Comparison with the spectrum calculated using an energy resolution of 1 meV. The splittings $\Delta$ and $\Delta^{\prime}$ are indicated in panel (c). Mode positions determined from the INS data are shown by pink circles (IN8) and triangles (ThALES); the orientation of the error bar indicates whether a point originates from a constant-$E$ or constant-$\bf Q$ scan. The grey curves show the dispersion calculated from linear spin-wave theory.
@@
From the near-tetragonal nature of the high-pressure phase, we assume that the SSM continues to provide a minimal model for the observed spin-wave dispersion. The physics of the AFM phase is that the stronger $J^{\prime}$ forces the spins to order ferromagnetically (FM) on the dimers and AFM between dimers, both in and out of plane [Fig. 1(c)]. The 2D AFM phase has two doubly degenerate spin-wave branches, a Goldstone mode and an optical mode. Although we have no dispersion measurements in the $c$ direction, the observed in-plane scattering intensity [Figs. 3(a)-(b)] motivates the introduction of an inter-layer interaction, $J_{c}$, between orthogonal dimers along the $(0,0,1)$ direction, as shown in Fig. 1(c). This model has the following properties: (i) the degeneracy of the low-energy mode is lifted, with a splitting $\Delta=8S\sqrt{J^{\prime}J_{c}}$ above the massless point; (ii) due to the AFM inter-layer stacking, the intensity of the Goldstone mode vanishes at ${\bf Q}=(0,1,0)$ and $(0,3,0)$, making the spectrum look gapped at these positions; (iii) there is an intensity asymmetry along $(0,q_{k},0)$ around $q_{k}=1$ and $3$, caused by the orthogonal arrangement of the dimers in the plane; (iv) the splitting between the low-energy and optical branches at the top of the former [e.g. ${\bf Q}=(0,1.5,0)$ and $(0.5,1,0)$] is $\Delta^{\prime}\approx 2SJ$.
@@
We refine the model parameters based on the experimentally determined mode positions and intensities with assistance from spinW Toth and Lake (2015), a software implementation of linear spin-wave theory, to obtain $J^{\prime}=4.28(14)$ meV, $J=2.36(26)$ meV (i.e. $J^{\prime}/J=1.8(2)$), and $J_{c}=0.053(3)$ meV. The predicted mode positions are shown in Fig. 2 as the orange symbols. The model spectra for an assumed instrumental energy resolution of 1 meV are shown in Figs. 3(c)-(d), and provide good agreement with the measurements in Figs. 3(a)-(b). In particular, the fitted spectrum confirms an AFM layer stacking, because FM stacking would result in Goldstone-mode intensity at $(0,1,0)$ and $(0,3,0)$, but not at $(1,1,0)$. We note that our deduced parameter values are not corrected for the quantum renormalization factor, $Z_{c}$, which to our knowledge has not been evaluated for the AFM phase of the SSM.
@@
Currently the most refined estimates for the interaction parameters of SrCu2(BO3)2 at ambient pressure are $J=81.5$ K (7.0 meV) and $J^{\prime}=51.3$ K (4.4 meV, $J^{\prime}/J=0.63$) Shi et al. (2022). While there is no exact model for the increase of $J^{\prime}/J$ with pressure, a linear decrease of both $J$ and $J^{\prime}$, with $\Delta J/P$ falling 3-4 times faster than $\Delta J^{\prime}/P$ Zayed et al. (2017), reproduces the $(H,P)$ phase diagram rather well in the dimer and plaquette phases Shi et al. (2022). Ab initio calculations predict a discontinuity in the parameters on entering the SSM-AFM phase Badrtdinov et al. (2020), and at higher pressures no information for $J$ or $J^{\prime}$ was available until now. Although it is unexpected that $J$ is suppressed so strongly, it is not implausible. Its size and sign are highly sensitive to the Cu-O-Cu bond angle following the Goodenough-Kanamori rules Goodenough (1955); Kanamori (1959), especially when this angle is close to 95 Mizuno et al. (1998). Measurements in SrCu2(BO3)2 [Figs. 1(a)-(b)] indicate a decrease from $98^{\circ}$ at ambient pressure to $94^{\circ}$ at 3.7 GPa Zayed (2010). The $J^{\prime}$ superexchange path is through a BO3 unit, which remains robust with pressure and locked in position by a mirror plane, becoming free to rotate out of the plane only beyond the monoclinic transition. Our results therefore quantify all of these considerations.
@@
Turning to $J_{c}$, we note first that the system is much softer along the ${\hat{c}}$ direction than in the plane, with the $c$ lattice parameter reduced by around 6% at 5.5 GPa, whereas $a$ and $b$ are reduced by less than 1% Loa et al. (2005); Haravifard et al. (2014). The buckling of the Cu-BO3 planes shown in Fig. 1(b) leads to two independent $J_{c}$ parameters Miyahara and Ueda (2003). However, the in-plane INS spectrum is sensitive only to their combined effect, and thus we fit a single effective $J_{c}$; further details are presented in Sec. S2A of the SM sup . In this quasi-2D system, the splitting $\Delta$ is a steep function of $J_{c}$ and in our INS experiment $\Delta=1.90(8)$ meV is measured to high accuracy, making our estimate of $J_{c}=0.053(3)$ meV at 5.5 GPa very precise. Although this small value reflects the absence of significant inter-layer superexchange paths, it is still thought that $J_{c}$ has stronger contributions from superexchange than from direct exchange, making its evolution under pressure difficult to predict. Nevertheless, the ambient-pressure $J_{c}$ is unlikely to be much larger than 0.05 meV. An indirect estimate of 0.7 meV, an order of magnitude above our result, was obtained as a 10% correction to the high-temperature magnetic susceptibility Miyahara and Ueda (2000, 2003), but was presumably a consequence of strong finite-size effects inherent to early numerical studies (since remedied by more modern technology Wietek et al. (2019)).
@@
Although a Heisenberg SSM with an inter-layer coupling captures all the measured features of the spin-wave dispersion in the high-pressure phase, we draw attention to two key points. First, SrCu2(BO3)2 has weak DM interactions at ambient pressure Miyahara and Ueda (2003); Kodama et al. (2005); Shi et al. (2022). These terms also act to gap the Goldstone mode, but the DM interaction required to create a gap of 1.9 meV in a 2D SSM is $0.7J^{\prime}$, as we show in Sec. 2B of the SM sup , making this a very unlikely mechanism to explain the INS spectrum. In our minimal model, we absolutely do not claim that DM interactions are no longer present, only that we cannot resolve them. In Sec. 2B we show also that even 10% DM terms in the lower-symmetry monoclinic phase of SrCu2(BO3)2 would provide only a minor correction to the Heisenberg interaction parameters. Second, we stress that linear spin-wave theory does not account for two-magnon scattering, possible two-magnon bound states, or any other effects of quantum fluctuations caused by the frustration remaining in the system, all of which, as we summarize in Sec. 2C of the SM sup , could contribute to the broad and weak neutron intensities observed above 5 meV in Figs. 2(d) and 3(a)-(b).
@@
The confirmation of a small $J_{c}$ is essential information for assessing the suitability of a 2D SSM to describe SrCu2(BO3)2. The fact that the dimer ($J$) bonds are forced into a FM configuration in the AFM phase makes the four $J_{c}$ bonds in an inter-layer tetrahedron fully unfrustrated [Fig. 1(c)], and hence they contribute an energy of order 1 K per site to the stability of a 3D AFM phase. In the dimer-singlet phase, by contrast, these bonds are simply ineffective. Concerning the plaquette phase, it was shown rather early that its very existence depends on having only a weak $J_{c}$ Koga (2000), and the propensity of $J_{c}$ to destabilize the plaquette phase has been used recently to extract a value $J_{c}\approx 0.2$ meV Vlaar and Corboz (2023). Our definitive $J_{c}$ can be used to refine the estimates of intra-layer energetics and their pressure evolution on which that estimate was based. Another delicate issue is whether the plaquette singlets form around the squares that contain the dimer bond (“full plaquette”) Zayed et al. (2017) or the squares that do not (“empty plaquette”) Corboz and Mila (2013). The minimal $J_{c}$ we consider should not affect plaquette formation Vlaar and Corboz (2023), but too little is known about the 3D plaquette order to make a statement on whether more complex inter-layer coupling could do so. Finally, because the plaquette-AFM transition of the 2D SSM has been proposed as a candidate deconfined quantum critical point (DQCP), and SrCu2(BO3)2 as a candidate material for its observation Lee et al. (2019); Yang et al. (2022); Cui et al. (2023), the effects of inter-layer coupling should definitely be included as part of evaluating such a scenario.
@@
Finally, we remark that our results do not provide an immediate understanding of the reported phase transitions at 8 K Guo et al. (2020) and 120 K Haravifard et al. (2014); Zayed et al. (2014); Guo et al. (2020), the parameters of our quasi-2D model being more likely to favor a single ordering transition at an intermediate temperature (of order $J^{\prime}/2$ Yasuda et al. (2005)). We performed detailed diffraction studies of the $(0,1,0)$ magnetic Bragg peak, as described in Sec. S1F of the SM sup , finding an unconventional thermal evolution with a characteristic temperature of 88 K. While one may speculate about the onset of 2D and of true 3D order, these contradictions underline the need for more systematic combined studies of the structure and magnetism at and above 3 GPa.